# IberFire - a detailed creation of a spatio-temporal dataset for wildfire risk assessment in Spain

### v1

@@ 2.10 Summary of external data sources
To ensure transparency and reproducibility, Table 4 provides direct links to the original raw datasets used in the construction of the IberFire datacube. These sources include official and publicly accessible repositories from European and national institutions. Each dataset listed in the table corresponds to one or more features included in the datacube, and their inclusion allows users to verify data provenance, explore updates, or perform additional processing tailored to specific needs.
@@
Auxiliary features Spain boundary Geometries https://simplemaps.com/gis/country/es
Fire history Historical fire data Geometries https://forest-fire.emergency.copernicus.eu/apps/data.request.form/
Geographical location Autonomous Communities Geometries https://www.arcgis.com/home/item.html?id=5f689357238847bc823a2fb164544a77
Land usage CLC_2006, CLC_2012, CLC_2018 100m https://land.copernicus.eu/en/products/corine-land-cover/
Topography Elevation, slope, aspect, roughness 30m https://portal.opentopography.org/raster?opentopoID=OTSDEM.032021.4326.3
Human activity Population density 1km https://hub.worldpop.org/geodata/listing?id=76
Distance to roads 100m https://hub.worldpop.org/geodata/summary?id=17504
Distance to waterways 100m https://hub.worldpop.org/geodata/summary?id=18002
@@
Meteorological 2m temperature, surface pressure, 10m u-wind, 10m v-wind 9km, hourly https://cds.climate.copernicus.eu/datasets/reanalysis-era5-land
2m temperature, 2m dewpoint temperature, precipitation, 10m u-wind, 10m v-wind 9km, daily https://cds.climate.copernicus.eu/datasets/derived-era5-land-daily-statistics
Vegetation FAPAR 1km, 10-daily https://land.copernicus.eu/en/products/vegetation/fraction-of-absorbed-photosynthetically-active-radiation-v2-0-1km
300m, 10-daily https://land.copernicus.eu/en/products/vegetation/fraction-of-absorbed-photosynthetically-active-radiation-v1-0-300m
LAI 1km, 10-daily https://land.copernicus.eu/en/products/vegetation/leaf-area-index-v2-0-1km
300m, 10-daily https://land.copernicus.eu/en/products/vegetation/leaf-area-index-300m-v1.0
NDVI 1km, 10-daily https://land.copernicus.eu/en/products/vegetation/normalised-difference-vegetation-index-v3-0-1km
300m, 10-daily https://land.copernicus.eu/en/products/vegetation/normalised-difference-vegetation-index-v2-0-300m
LST 9km, daily https://cds.climate.copernicus.eu/datasets/derived-era5-land-daily-statistics?tab=overview
5km, hourly https://land.copernicus.eu/en/products/temperature-and-reflectance/hourly-land-surface-temperature-global-v1-0-5km
5km, hourly https://land.copernicus.eu/en/products/temperature-and-reflectance/hourly-land-surface-temperature-global-v2-0-5km
SWI_01, SWI_05, SWI_10, SWI_20 12.5km, daily https://land.copernicus.eu/en/products/soil-moisture/daily-soil-water-index-global-v3-0-12-5km
@@ 3 Data Records
The IberFire datacube comprises approximately $6.8\cdot 10^{9}$ individual cells, resulting from the combination of $1188$ values along the $x$ coordinate, 920 along the $y$ coordinate, and $6241$ distinct timestamps. Each $(x,y,t)$ cell stores a vector of 260 features that include static and dynamic variables relevant to wildfire risk modelling. At any given cell, all 260 feature values can be retrieved for analytical or predictive purposes. This chapter presents a detailed description of the 260 features, organised into structured tables that replicate the format found on IberFire when opened with xarray.
The datacube is publicly available on Zenodo (https://zenodo.org/records/15225886), and the variables are organised in the following tables according to the categories defined in Chapter 2. Table 5 presents the three auxiliary features, two fire-related variables, and five geographic context features. Table 6 details the 63 variables derived from the 2006 Corine Land Cover dataset. Table 7 then presents the 15 topographical and 21 human activity features. Lastly, Table 8 summarises the 17 meteorological variables and 8 vegetation indices included in the dataset. IberFire contains 126 additional features derived from the 2012 and 2018 CLC datasets. As those features follow the same structure as the 2006 version, they are not described separately.
@@
x_index X-coordinate index values Integer in [0, 1187]
y_index Y-coordinate index values Integer in [0, 919]
is_spain Binary mask indicating the Spanish region 0 (outside Spain), 1 (inside Spain)
@@
x_coordinate X-coordinate values in the EPSG:3035 reference system Meters
y_coordinate Y-coordinate values in the EPSG:3035 reference system Meters
is_sea Binary indicator denoting whether a cell lies over the open sea. The data was derived from the Copernicus Global Digital Elevation Model 0 (land), 1 (sea)
is_waterbody Binary indicator denoting whether a cell lies over inland water (e.g., lakes, rivers). The data was derived from the Copernicus Global Digital Elevation Model 0 (non-water), 1 (inland water)
AutonomusCommunity The Autonomous Communities code Label from 00 to 19
@@
CLC_2006_1 Proportion of original grid cells labelled as class 1 in CLC in the $1km\times 1km$ cell
CLC_2006_44 Proportion of original grid cells labelled as class 44 in CLC in the $1km\times 1km$ cell
CLC_2006_urban_fabric_proportion Proportion of original cells labelled as class 1 or 2 in CLC in the $1km\times 1km$ cell
CLC_2006_industrial_proportion Proportion of original cells labelled as class 3, 4, 5 or 6 in CLC in the $1km\times 1km$ cell
CLC_2006_mine_proportion Proportion of original cells labelled as class 7, 8 or 9 in CLC in the $1km\times 1km$ cell
CLC_2006_artificial _vegetation_proportion Proportion of original cells labelled as class 10 or 11 in CLC in the $1km\times 1km$ cell
CLC_2006_arable_land_proportion Proportion of original cells labelled as class 12, 13 or 14 in CLC in the $1km\times 1km$ cell
CLC_2006_permanent_crops_proportion Proportion of original cells labelled as class 15, 16 or 17 in CLC in the $1km\times 1km$ cell
CLC_2006_heterogeneous _agricultural_proportion Proportion of original cells labelled as class 19, 20 or 21 in CLC in the $1km\times 1km$ cell
CLC_2006_forest_proportion Proportion of original cells labelled as class 23, 24 or 25 in CLC in the $1km\times 1km$ cell
CLC_2006_scrub_proportion Proportion of original cells labelled as class 26, 27, 28 or 29 in CLC in the $1km\times 1km$ cell
CLC_2006_open_space_proportion Proportion of original cells labelled as class 30, 31, 32, 33 or 34 in CLC in the $1km\times 1km$ cell
CLC_2006_inland_wetlands_proportion Proportion of original cells labelled as class 35 or 36 in CLC in the $1km\times 1km$ cell
CLC_2006_maritime_wetlands_proportion Proportion of original cells labelled as class 37, 38 or 39 in CLC in the $1km\times 1km$ cell
CLC_2006_inland_waters_proportion Proportion of original cells labelled as class 40 or 41 in CLC in the $1km\times 1km$ cell
CLC_2006_marine_waters_proportion Proportion of original cells labelled as class 42, 43 or 44 in CLC in the $1km\times 1km$ cell
CLC_2006_artificial_proportion Proportion of original cells labelled as classes from 1 to 11 in CLC in the $1km\times 1km$ cell
CLC_2006_agricultural_proportion Proportion of original cells labelled as classes from 12 to 22 in CLC in the $1km\times 1km$ cell
CLC_2006_forest_and_semi _natural_proportion Proportion of original cells labelled as classes from 23 to 34 in CLC in the $1km\times 1km$ cell
CLC_2006_wetlands_proportion Proportion of original cells labelled as classes from 35 to 39 in CLC in the $1km\times 1km$ cell
CLC_2006_waterbody_proportion Proportion of original cells labelled as classes from 40 to 44 in CLC in the $1km\times 1km$ cell
Table 6: Corine Land Cover features (corresponding to 2006). All features correspond to proportions, with values ranging from 00 to $1$. The top part describes the 44 raw classes of the CLC dataset. The middle part corresponds to the 14 intermediate aggregation levels. The lower part of the table represents the 5 higher clustering levels of CLC. This hierarchical ordering is defined in Table 2.
@@
elevation_mean Mean elevation in the 1 km $\times$ 1 km grid cell. The elevation is computed from the Digital Elevation Model (DEM) at 30 m resolution Meters
elevation_stdev Standard deviation in the 1 km $\times$ 1 km grid cell. The elevation is computed from the Digital Elevation Model (DEM) at 30 m resolution Meters
slope_mean Mean slope in the 1 km $\times$ 1 km grid cell. The slope is computed from the Digital Elevation Model (DEM) at 30 m resolution -
slope_stdev Standard deviation slope in the 1 km $\times$ 1 km grid cell. The slope is computed from the Digital Elevation Model (DEM) at 30 m resolution -
roughness_mean Mean roughness in the 1 km $\times$ 1 km grid cell. The roughness is computed from the Digital Elevation Model (DEM) at 30 m resolution -
rougness_stdev Standard deviation roughness in the 1 km $\times$ 1 km grid cell. The roughness is computed from the Digital Elevation Model (DEM) at 30 m resolution -
aspect_1 Proportion of the aspect class 1 in the 1 km $\times$ 1 km grid cell. The aspect classes are computed from the digital elevation model (DEM) and represent the orientation of the terrain Proportion [0,1]
aspect_8 Proportion of the aspect class 8 in the 1 km $\times$ 1 km grid cell. The aspect classes are computed from the digital elevation model (DEM) and represent the orientation of the terrain Proportion [0,1]
aspect_NODATA Proportion of the aspect class NODATA in the 1 km $\times$ 1 km grid cell. The aspect classes are computed from the digital elevation model (DEM) and represent the orientation of the terrain Proportion [0,1]
dist_to_roads_mean Mean distance to roads in the 1 km $\times$ 1 km grid cell Kilometers
dist_to_roads_stdev Standard deviation of the distance to roads in the 1 km $\times$ 1 km grid cell Kilometers
dist_to_waterways_mean Mean distance to waterways in the 1 km $\times$ 1 km grid cell Kilometers
dist_to_waterways_stdev Standard deviation of the distance to waterways in the 1 km $\times$ 1 km grid cell Kilometers
dist_to_railways_mean Mean distance to railways in the 1 km $\times$ 1 km grid cell -
dist_to_railways_stdev Standard deviation of the distance to railways in the 1 km $\times$ 1 km grid cell -
is_holiday Binary mask indicating whether it is a holiday in the 1 km $\times$ 1 km grid cell in that time or not 0 (working day), 1 (holiday)
is_natura2000 Binary mask indicating whether the 1 km $\times$ 1 km grid cell is part of the Natura 2000 network or not 0 (is not in Natura 2000), 1 (is in Natura 2000)
popdens_2008 Mean population density in the 1 km $\times$ 1 km grid cell for the year 2008 People/$km^{2}$
popdens_2020 Mean population density in the 1 km $\times$ 1 km grid cell for the year 2020 People/$km^{2}$
@@
t2m_mean The mean temperature of air measured at 2m above the surface of the land, sea or inland waters Celsius
t2m_min The minimum temperature of air measured at 2m above the surface of the land, sea or inland waters Celsius
t2m_max The maximum temperature of air measured at 2m above the surface of the land, sea or inland waters Celsius
t2m_range The range temperature of air measured at 2m above the surface of the land, sea or inland waters Celsius
RH_mean The mean relative humidity of air [0, 100] in %
RH_min The minimum relative humidity of air [0, 100] in %
RH_max The maximum relative humidity of air [0, 100] in %
RH_range The range relative humidity of air [0, 100] in %
surface_pressure_mean The mean surface pressure of air Hectopascal
surface_pressure_min The minimum surface pressure of air Hectopascal
surface_pressure_max The maximum surface pressure of air Hectopascal
surface_pressure_range The range surface pressure of air Hectopascal
total_precipitation _mean Mean of the hourly values of the total precipitation variable from ERA5-Land millimetres (1l/$m^{2}$)
wind_speed_mean The mean wind speed derived from the hourly u-component and v-component of wind Meters per second (m/s)
wind_speed_max The maximum wind speed derived from the hourly u-component and v-component of wind Meters per second (m/s)
wind_direction_mean The mean wind direction (where the wind comes) derived from the mean u-component and v-component of wind Degrees
wind_direction_at _max_speed The wind direction (where the wind comes) where the maximum wind speed happened. Derived after calculating the hourly wind speed from the hourly u-component and v-component of wind Degrees
fapar Fraction of Absorbed Photosynthetically Active Radiation, the fraction of the solar radiation absorbed by live plants for photosynthesis -
lai Leaf Area Index, representing the half of the total green canopy area per unit horizontal ground area -
LST Land Surface Temperature is the temperature of the surface of the Earth Kelvin
ndvi Normalized Difference Vegetation Index, an indicator of the greenness of the biomes -
SWI_001 Soil Water Index at T=1, provides daily updates on the moisture conditions in different soil layers [0, 100] in %
SWI_005 Soil Water Index at T=5, provides daily updates on the moisture conditions in different soil layers [0, 100] in %
SWI_010 Soil Water Index at T=10, provides daily updates on the moisture conditions in different soil layers [0, 100] in %
SWI_020 Soil Water Index at T=20, provides daily updates on the moisture conditions in different soil layers [0, 100] in %

### v2

@@ 2.10 Summary of external data sources
To ensure transparency and reproducibility, Table 5 provides direct links to the original raw datasets used in the construction of the IberFire datacube. Each dataset listed in the table corresponds to one or more features included in the datacube. These sources include official and publicly accessible repositories from European and national institutions, and their inclusion allows users to verify data provenance or perform additional processing tailored to specific use cases, like model deployment.
@@
Auxiliary features Spain boundary Geometries https://simplemaps.com/gis/country/es [simplemaps-spain]
Fire history Historical fire data Geometries https://forest-fire.emergency.copernicus.eu/apps/data.request.form/ [effis]
FWI 27.5km https://ewds.climate.copernicus.eu/datasets/cems-fire-historical-v1?tab=overview [FWI_data]
Geographical location Autonomous Communities Geometries https://www.arcgis.com/home/item.html?id=5f689357238847bc823a2fb164544a77 [area2018limites]
Land usage CLC_2006, CLC_2012, CLC_2018 100m https://land.copernicus.eu/en/products/corine-land-cover [CLC_2006_version] [CLC_2012_version] [CLC_2018_version]
Topography Elevation, slope, aspect, roughness 30m https://portal.opentopography.org/raster?opentopoID=OTSDEM.032021.4326.3 [EU_DEM]
Human activity Population density 1km https://hub.worldpop.org/doi/10.5258/SOTON/WP00674 [popdens]
Distance to roads 100m https://hub.worldpop.org/geodata/summary?id=17504 [distance_to_roads]
Distance to waterways 100m https://hub.worldpop.org/geodata/summary?id=18002 [distance_to_waterways]
@@
Meteorological 2m temperature, 2m dewpoint temperature, precipitations, 10m u-wind, 10m v-wind 9km, hourly https://cds.climate.copernicus.eu/datasets/reanalysis-era5-land?tab=overview [era5_hourly_data]
2m temperature, surface pressure, 10m u-wind, 10m v-wind 9km, daily https://cds.climate.copernicus.eu/datasets/derived-era5-land-daily-statistics?tab=overview [era5_postprocessed_data]
Vegetation FAPAR 1km, 10-daily https://land.copernicus.eu/en/products/vegetation/fraction-of-absorbed-photosynthetically-active-radiation-v2-0-1km [FAPAR_v1]
300m, 10-daily https://land.copernicus.eu/en/products/vegetation/fraction-of-absorbed-photosynthetically-active-radiation-v1-0-300m [FAPAR_v2]
LAI 1km, 10-daily https://land.copernicus.eu/en/products/vegetation/leaf-area-index-v2-0-1km [LAI_v1]
300m, 10-daily https://land.copernicus.eu/en/products/vegetation/leaf-area-index-300m-v1.0 [LAI_v2]
NDVI 1km, 10-daily https://land.copernicus.eu/en/products/vegetation/normalised-difference-vegetation-index-v3-0-1km [NDVI_v1]
300m, 10-daily https://land.copernicus.eu/en/products/vegetation/normalised-difference-vegetation-index-v2-0-300m [NDVI_v2]
LST 9km, daily https://cds.climate.copernicus.eu/datasets/derived-era5-land-daily-statistics?tab=overview [era5_postprocessed_data]
5km, hourly https://land.copernicus.eu/en/products/temperature-and-reflectance/hourly-land-surface-temperature-global-v1-0-5km [LST_v2]
5km, hourly https://land.copernicus.eu/en/products/temperature-and-reflectance/hourly-land-surface-temperature-global-v2-0-5km [LST_v3]
SWI_001, SWI_005, SWI_010, SWI_020 12.5km, daily https://land.copernicus.eu/en/products/soil-moisture/daily-soil-water-index-global-12-5km [SWI_dataset]
@@ 3 Data Records
IberFire comprises approximately $6.8\times 10^{9}$ individual cells, resulting from the combination of $1188$ values along the $x$ coordinate, 920 along the $y$ coordinate, and $6241$ distinct timestamps. When IberFire is opened with the xarray Python package, 261 ‘data variables’ are shown. These are all the spatio-temporal and spatial-only features that can be retrieved for each cell, including the low update frequency features stored as spatial-only features.
Therefore, the three versions of the CLC dataset, each with 63 features, along with the 13 yearly population density features, result in a total of 261 ‘data variables’. However, when the correct versions of CLC and population density are selected for each cell, there are a total of $261-63\cdot 2-12=123$ different ‘real’ variables. From those 123 features, the auxiliary features are not intended to be used for modelling purpose, but for data manipulation. Therefore, IberFire provides a total of 120 different features for modelling.
Since opening the dataset initially displays the ‘data variables’, this section presents a detailed description of them, organised into structured tables that replicate the format found on IberFire when opened with xarray.
The datacube is publicly available on Zenodo (https://zenodo.org/records/15798999), and the variables are organised in the following tables according to the categories defined in Section 2. Table 6 presents the three auxiliary features, two fire-related variables, and five geographic context features. Table 7 details the 63 variables derived from the 2006 CLC dataset (the 2012 and 2018 CLC version are not described since they mirror the 2006 version). Table 8 then presents the 15 topographical and 21 human activity features (including the repeated population density features). Lastly, Table 9 summarises the 17 meteorological variables and 8 vegetation indices included in the dataset.
@@
x_index X-coordinate index values. Integer in [0, 1187]
y_index Y-coordinate index values. Integer in [0, 919]
is_spain Binary mask indicating the Spanish region. 0 (outside Spain), 1 (inside Spain)
@@
x_coordinate X-coordinate values in the EPSG:3035 reference system. Metres (float)
y_coordinate Y-coordinate values in the EPSG:3035 reference system. Metres (float)
is_sea Binary indicator denoting whether a cell lies over the open sea. 0 (land), 1 (sea)
is_waterbody Binary indicator denoting whether a cell lies over inland water (e.g., lakes, rivers). 0 (non-water), 1 (inland water)
AutonomusCommunity The Autonomous Communities code. Label from 00 to 19
@@
Label 3 - Raw CLC Classes (1 - 44)
CLC_2006_1 Proportion of the original 100m $\times$ 100m cells labelled as CLC class 1 in the 1km $\times$ 1km cell
CLC_2006_44 Proportion of the original 100m $\times$ 100m cells labelled as CLC class 44 in the 1km $\times$ 1km cell
Label 2 - Intermediate aggregations
CLC_2006_urban_fabric_proportion Sum of CLC_2006_1 - CLC_2006_2
CLC_2006_industrial_proportion Sum of CLC_2006_3 - CLC_2006_6
CLC_2006_mine_proportion Sum of CLC_2006_7 - CLC_2006_9
CLC_2006_artificial_vegetation_proportion Sum of CLC_2006_10 - CLC_2006_11
CLC_2006_arable_land_proportion Sum of CLC_2006_12 - CLC_2006_14
CLC_2006_permanent_crops_proportion Sum of CLC_2006_15 - CLC_2006_17
CLC_2006_heterogeneous_agriculture_proportion Sum of CLC_2006_19 - CLC_2006_22
CLC_2006_forest_proportion Sum of CLC_2006_23 - CLC_2006_25
CLC_2006_scrub_proportion Sum of CLC_2006_26 - CLC_2006_29
CLC_2006_open_space_proportion Sum of CLC_2006_30 - CLC_2006_34
CLC_2006_inland_wetlands_proportion Sum of CLC_2006_35 - CLC_2006_36
CLC_2006_maritime_wetlands_proportion Sum of CLC_2006_37 - CLC_2006_39
CLC_2006_inland_waters_proportion Sum of CLC_2006_40 - CLC_2006_41
CLC_2006_marine_waters_proportion Sum of CLC_2006_42 - CLC_2006_44
Label 1 - High level aggregations
CLC_2006_artificial_proportion Sum of CLC_2006_1 - CLC_2006_11
CLC_2006_agricultural_proportion Sum of CLC_2006_12 - CLC_2006_22
CLC_2006_forest_and_semi_natural_proportion Sum of CLC_2006_23 - CLC_2006_34
CLC_2006_wetlands_proportion Sum of CLC_2006_35 - CLC_2006_39
CLC_2006_waterbody_proportion Sum of CLC_2006_40 - CLC_2006_44
Table 7: Corine Land Cover features (corresponding to 2006). All features correspond to proportions, with values ranging from 0 to 11. The top part describes the 44 raw classes of the CLC dataset. The middle part corresponds to the 14 intermediate aggregation levels. The lower part of the table represents the 5 higher clustering levels of CLC. This hierarchical ordering is defined in Table 3.
@@
elevation_mean Mean elevation in the 1 km $\times$ 1 km grid cell. Metres (float)
elevation_stdev Standard deviation of elevation in the 1 km $\times$ 1 km grid cell. Metres (float)
slope_mean Mean slope in the 1 km $\times$ 1 km grid cell. -
slope_stdev Standard deviation slope in the 1 km $\times$ 1 km grid cell. -
roughness_mean Mean roughness in the 1 km $\times$ 1 km grid cell. -
rougness_stdev Standard deviation roughness in the 1 km $\times$ 1 km grid cell. -
aspect_1 Proportion of the aspect class 1 in the 1 km $\times$ 1 km grid cell. Proportion [0,1]
aspect_8 Proportion of the aspect class 8 in the 1 km $\times$ 1 km grid cell. Proportion [0,1]
aspect_NODATA Proportion of the aspect class NODATA in the 1 km $\times$ 1 km grid cell. Proportion [0,1]
dist_to_roads_mean Mean distance to roads in the 1 km $\times$ 1 km grid cell. Kilometres (float)
dist_to_roads_stdev Standard deviation of the distance to roads in the 1 km $\times$ 1 km grid cell. Kilometres (float)
dist_to_waterways_mean Mean distance to waterways in the 1 km $\times$ 1 km grid cell. Kilometres (float)
dist_to_waterways_stdev Standard deviation of the distance to waterways in the 1 km $\times$ 1 km grid cell. Kilometres (float)
dist_to_railways_mean Mean distance to railways in the 1 km $\times$ 1 km grid cell. -
dist_to_railways_stdev Standard deviation of the distance to railways in the 1 km $\times$ 1 km grid cell. -
is_holiday Binary mask indicating whether it is a holiday in the 1 km $\times$ 1 km grid cell in that time or not. 0 (working day), 1 (holiday)
is_natura2000 Binary mask indicating whether the 1 km $\times$ 1 km grid cell is part of the Natura 2000 network or not. 0 (is not in), 1 (is in)
[…]
@@
t2m_mean The mean temperature of air measured at 2m above the surface of the land, sea or inland waters. Degrees Celsius
t2m_min The minimum temperature of air measured at 2m above the surface of the land, sea or inland waters. Degrees Celsius
t2m_max The maximum temperature of air measured at 2m above the surface of the land, sea or inland waters. Degrees Celsius
t2m_range The range temperature of air measured at 2m above the surface of the land, sea or inland waters. Degrees Celsius
RH_mean The mean relative humidity of air. [0, 100] in %
RH_min The minimum relative humidity of air. [0, 100] in %
RH_max The maximum relative humidity of air. [0, 100] in %
RH_range The range relative humidity of air. [0, 100] in %
surface_pressure_mean The mean surface pressure of air. Hectopascal
surface_pressure_min The minimum surface pressure of air. Hectopascal
surface_pressure_max The maximum surface pressure of air. Hectopascal
surface_pressure_range The range surface pressure of air. Hectopascal
[…]
wind_speed_mean The mean wind speed derived from the hourly u-component and v-component of wind. Metres per second (m/s)
wind_speed_max The maximum wind speed. Metres per second (m/s)
wind_direction_mean The mean wind direction (where the wind comes). Degrees
wind_direction_at _max_speed The wind direction (where the wind comes) where the maximum wind speed happened. Degrees
FAPAR Fraction of Absorbed Photosynthetically Active Radiation, the fraction of the solar radiation absorbed by live plants for photosynthesis. -
LAI Leaf Area Index, representing the half of the total green canopy area per unit horizontal ground area. -
LST Land Surface Temperature is the temperature of the surface of the Earth. Degrees Kelvin
ndvi Normalised Difference Vegetation Index, an indicator of the greenness of the biomes. -
SWI_001 Soil Water Index at T=1, the moisture humidity conditions of the soil. [0, 100] in %
SWI_005 Soil Water Index at T=5, the moisture humidity conditions of the soil. [0, 100] in %
SWI_010 Soil Water Index at T=10, the moisture humidity conditions of the soil. [0, 100] in %
SWI_020 Soil Water Index at T=20, the moisture humidity conditions of the soil. [0, 100] in %